A capacitor has a capacitance of 33.68 pF. What potential difference would be required to store 69.82 pC?


1
Expert's answer
2021-05-02T18:31:22-0400

To be given in question

Charge(Q)=69.82PC ="69.82\times10^{-12} C"

Capacity(C)=33.68PF ="33.68\times10^{-12}F"

To be asked in question

Potential difference (∆v)=?

We know that

"∆V=\frac{Q}{C}\rightarrow(1)"

eqution (2)put value

"∆V=\frac{69.82\times 10^{-12}}{33.68\times10^{-12}} V"

"\u2206V=2.07304V"